A cylinder has a radius of 4 meters and an approximate volume of 351.9. To the nearest whole meter, what is the height of this cylinder?
Answer by Boreal(15235)

V=pi*r^2*h
351.9=pi*16*h
h=351.9/(16pi)=7.00 meters or 7 meters